OpenStudy (anonymous):

Differentiate the function g(x) = ln [x*sqrt(x^2 - 1)]

OpenStudy (anonymous):

\[g(x) = \ln (x \sqrt(x^2 - 1))\] this one?

OpenStudy (anonymous):

use chain rule

OpenStudy (anonymous):

start with \[\ln(x)+\frac{1}{2}\ln(x+1)+\frac{1}{2}\ln(x-1)\] and it will be easier

OpenStudy (anonymous):

sooo g'(x) = 1/x + 1/x(x+1) + 1/x(x-1) ?

OpenStudy (anonymous):

stupid i know but i'm having trouble simplifying that what do i do

OpenStudy (anonymous):

g'(x) = 1/x + 1/2(1/(x+1)) + 1/2(1/(x-1)) = 1/x + 1/2(x+1) + 1/2(x-1)

OpenStudy (anonymous):

well yeah i got that sorry i typed it wrong but i need to simplify that